OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 2 red marbles, 5 white marbles, and 3 blue marbles. You return the first marble to the bag and then choose again. Find P(red and blue).

OpenStudy (turingtest):

(red+blue)/total=5/10=1/2